Estimate the amount of methane that can be recovered from one ton of municipal solid waste(MSW)

moisture content in MSW =20%

Mass retention time =30d

Volatile solids(VS)=0.93 * total solids(TS)

Biodegradable Volatile Solids,BVS=0.7*VS

Expected BVS conversion efficiency=95%

Gas production=8ft^3/lb BVS destroyed

Energy content of biogas=500Btu/ft^3

Conversion Factor:10^5BTU/therm

Monetary Value:$0.6/Therm
Answer #1

Answer = total value of the biogas produced

● Determining the mass of VS in 1 ton of of MSW -:    VS = 1.0 ton × 1000 kg/ton × 0.80 × 0.93 = 744 kg

●Determining the mass of BVS destroyed -:
Mass of BVS destroyed = 744 kg × 0.70 × 0.95 = 494.8 kg

● Determinig the volume of biogas produced -:
Biogas produced = 494.8 kg × 8 ft3 /lb = 3960 ft3

● Determining total energy value of the biogas-:
Energy value = 3960 ft3 × 500 Btu/ft3 = 2 × 106Btu

● Determining total value of the biogas produced -:

Value = 2 × 106 Btu × 0.6/Therm = 12 × 105 USD/ton

Hope you have got your answer. Give POSITIVE FEEDBACK by clicking on thumbs up so we can serve you better.
THANK YOU.